### 18 Jan 2024 v1

# Code Prompting Elicits Conditional Reasoning Abilities in Text+Code LLMs

Abstract
Reasoning is a fundamental component for achieving language understanding. Among the multiple types of reasoning, conditional reasoning, the ability to draw different conclusions depending on some condition, has been understudied in large language models (LLMs). Recent prompting methods, such as chain of thought, have significantly improved LLMs on reasoning tasks. Nevertheless, there is still little understanding of what triggers reasoning abilities in LLMs. We hypothesize that code prompts can trigger conditional reasoning in LLMs trained on text and code. We propose a chain of prompts that transforms a natural language problem into code and prompts the LLM with the generated code. Our experiments find that code prompts exhibit a performance boost between $2.6$ and $7.7$ points on GPT 3.5 across multiple datasets requiring conditional reasoning. We then conduct experiments to discover how code prompts elicit conditional reasoning abilities and through which features. We observe that prompts need to contain natural language text accompanied by high-quality code that closely represents the semantics of the instance text. Furthermore, we show that code prompts are more efficient, requiring fewer demonstrations, and that they trigger superior state tracking of variables or key entities. [^1] Code, prompt templates, prompts, and outputs are publicly available at https://github.com/UKPLab/arxiv2024-conditional-reasoning-llms.
Code Prompting Elicits Conditional Reasoning Abilities
in Text+Code LLMs
[…]
1 Introduction
Reasoning is a fundamental component of both human and artificial intelligence (AI). Some reasoning types have received much attention in the form of new benchmarks and models from the AI community in the last few years, such as mathematical (Patel et al., 2021; Chen et al., 2021; Cobbe et al., 2021), logical (Liu et al., 2020, 2023a; Sinha et al., 2019), and commonsense reasoning (Madaan et al., 2022; Liu et al., 2022a, b; Wang et al., 2023). However, other important reasoning types, such as conditional reasoning, remain understudied. Conditional reasoning draws alternative conclusions depending on the fulfillment of certain conditions. It is also a fundamental form of logical reasoning useful in many scenarios, such as in chatbots (e.g., ChatGPT), to answer real-world questions like eligibility for a visa or a loan. Despite the recent introduction of some benchmarks (Saeidi et al., 2018; Sun et al., 2022; Kazemi et al., 2023), conditional reasoning abilities of LLMs remain unknown.
Figure 1: Code prompting converts a natural language problem into a code prompt and prompts a large language model with such code to generate a natural language answer.
Recently, researchers have improved performance on reasoning tasks by combining LLMs with symbolic interpreters such as the Python runtime  (Gao et al., 2023; Chen et al., 2023; Lyu et al., 2023) or SATisfiability solvers  (Ye et al., 2023). Here, LLMs pretrained on code or a combination of text and code (henceforth text+code LLM) are used to convert the input task into a symbolic language (e.g., Python or SAT problems), which is then fed into an external interpreter to make use of its symbolic execution. However, offloading the final reasoning task to an external execution module does not help with, or sometimes confounds, our understanding of the reasoning abilities of the model because part of the performance gain may come from the correct execution of the external interpreter. In particular, the fundamental questions of what contributes to the reasoning abilities and how reasoning abilities are triggered are still open questions needing further understanding. Nevertheless, pretraining on code is considered an important component that contributes to and explains the improved reasoning ability of LLMs. State-of-the-art LLMs such as GPT 3.5 (Kojima et al., 2022), GPT 4 (OpenAI, 2023), PaLM (Chowdhery et al., 2023), and Mistral 7B (Jiang et al., 2023) have been pretrained on both text and code and have demonstrated considerable boosts in many reasoning benchmarks.
In this work, we investigate which aspects of code prompts, and in which way, elicit conditional reasoning abilities of GPT 3.5, a text+code LLM, across two datasets: (1) ConditionalQA (Sun et al., 2022), a scenario-based question answering (QA) dataset and (2) BoardgameQA (Kazemi et al., 2023), a boardgame-based QA dataset with conflicting rules. To understand the true benefit of code as an intermediate representation, we devise a chain of prompts, code prompting, that transform a natural language (NL) task into code and prompt the LLM with the generated code to elicit reasoning abilities. An illustration is provided in Figure 1. This setup has multiple benefits. Firstly, we fully utilize LLM’s reasoning ability without offloading any subtask to an external interpreter. In this way, we can make a fair comparison between text and code prompts without external variables such as interpreters. Secondly, enforcing compilability and executability of the generated code is not required, resulting in fewer interpreter-driven errors.
Our contributions are summarized as follows:
  • We devise a chain of prompts that transforms a NL task into code to trigger conditional reasoning abilities in text+code LLMs.
  • We conduct a comprehensive study to compare code prompts with text prompts, showing i) large performance gains between $2.6$ and $7.7$ points while ii) being more efficient w.r.t. the number of demonstrations.
  • We conduct an extensive analysis of why code prompts efficiently elicit conditional reasoning abilities in text+code LLMs, showing that prompting with code results in largely improved variable state tracking.
Figure 2: Code Prompts converts a natural language problem instance into code before solving it with a large language model.
2 Related Work
Most works generating code to solve natural language tasks use an external symbolic interpreter to run such code. Chen et al. (2023) and Gao et al. (2023) showed consistent gains on mathematical problems, symbolic reasoning, and algorithmic problems by using LLMs aided by external code interpreters. Lyu et al. (2023) show further gains in multi-hop QA, planning, and relational inference. Instead of leveraging a code interpreter, Ye et al. (2023) used an automated theorem prover with declarative code and showed consistent gains w.r.t. imperative code-interpreter-aided LLMs on arithmetic reasoning, logical reasoning, symbolic reasoning, and regex synthesis tasks. Lastly, Pan et al. (2023) did not use any interpreter and instead created programs composed of multiple subroutines and used smaller specialized models to run them. In this way, they outperform text prompts on text LLMs for fact-checking tasks.
All these works employ an external solver system to run the code; therefore, the LLM is not conducting part of the reasoning. However, some works (Madaan et al., 2022; Liu et al., 2023b) suggest that LLMs of code may possess superior reasoning abilities than LLMs of text (i.e., pretrained only on natural language text). Madaan et al. (2022) observed improved commonsense reasoning, while Liu et al. (2023b) showed superior causal reasoning. This last work is based on the intuition that the large amount of if statements in the pretraining corpus enhances the causal reasoning abilities because they represent explicit causal relations. They conducted experiments on abductive and counterfactual reasoning tasks and showed that translating NL problems into code and then generating functions that return the answer to the problem with a code LLM outperforms prompting the same NL task in a text LLM. However, most popular LLMs are trained on text and code (e.g., GPT 3.5; Kojima et al. 2022, GPT 4; OpenAI 2023) and it remains unknown whether the reported performance gains come from the prompt format, the abilities of the LLM, or a combination of both. Therefore, it remains unclear whether code prompts also outperform text prompts in text+code models and, if so, why code prompts elicit reasoning abilities. In our work, we aim to answer all these questions.
As far as we know, only the concurrent work of (Hussain et al., 2023) investigates the conditional reasoning abilities of LLMs. However, they only analyze the abilities of text-only LLMs after training them on ConditionalQA (Sun et al., 2022). Instead, we investigate how to use prompts to elicit conditional reasoning in text+code LLMs.
3 Code Prompting
3.1 Definition
We hypothesize that querying text+code LLMs with instances translated to (pseudo)-code will result in improved conditional reasoning capabilities. Our hypothesis is motivated by previous results showing that program-aided LMs exhibit superior results on reasoning tasks than regular text prompts (Gao et al., 2023; Chen et al., 2023; Lyu et al., 2023; Ye et al., 2023). Thus, it comes naturally that prompting text+code models with instances translated to code could cause the underlying model to exhibit superior reasoning abilities when compared to text prompts. We formally define our hypothesis as follows:
$\sum_{p\in P}\sigma(LLM(T(p)),p)\leq\sum_{p\in P}\sigma(LLM(C(p)),p)$
where $p$ is a problem instance from the set of problems $P$, $\sigma$ is an evaluation function that returns the quality of an answer for a given problem, and $T$ and $C$ are functions that create a natural language (Text) prompt and Code prompt, respectively, for the same given problem.
We define code prompts as prompts that model a natural language (NL) problem with code. The code contains the logical structure needed to solve the problem, along with the original natural language text as code comments. To solve an NL task with code prompts, we define a chain of prompts that i) transforms the NL text into code and ii) prompts using this code to generate the answer in natural language. Figure 1 illustrates this pipeline.
The prompt that transforms the NL problem to code is comprised of code that closely follows the original NL text. In particular, it creates variables for the key entities in the question and documents and if blocks for the documents, representing their conditional statements. Figure 2 illustrates this transformation.
3.2 Coding Features
To generate code as close as possible to the NL text, we use a programming language based on a simplification of Python. We only use boolean variables or variables that contain lists of strings. Variables follow the snake case naming convention. We also employ if statements to model conditional reasoning, but we do not use loops, functions, or classes. We create a code comment with the original NL text for each input sentence, and right after the code comment, we generate the code that represents the semantics of that sentence. However, we do not enforce that the generated code should be a runnable Python script.
4 Experimental Setup
4.1 Datasets
Throughout our experiments, we focus on two conditional question-answering QA datasets: ConditionalQA (CondQA; Sun et al., 2022) and BoardgameQA (BGQA; Kazemi et al., 2023). Solving these datasets requires advanced conditional and compositional reasoning capabilities.
ConditionalQA is a QA dataset where the answers are applicable under specific scenarios (i.e., conditional answers). Therefore, along with each question, the dataset provides a scenario that describes the background of the person posing such a question. Questions require multi-hop, compositional, and conditional logic over documents about public policies (e.g., the eligibility for a subsidy). Answers can be a span of the document, yes, and no. We use an oracle retriever to select the relevant passages to the question so that we can isolate the analysis of conditional reasoning abilities in LLMs from the retrieval component. The expected output is a chain of thought (CoT; Wei et al. 2022) followed by the final answer. To create the CoT, we use the annotated evidence sentences.
BoardgameQA is a dataset that evaluates the ability to reason with contradictory information guided by preferences. For example, given a question about traveling abroad, information found online about regulations can be contradictory because rules may change over time. Answering questions in this dataset requires complex multi-hop reasoning with conditional, deductive, and compositional abilities. The domain of the problems is board games, which allows us to analyze the conditional reasoning abilities in a completely different domain from CondQA. BGQA is divided into multiple partitions focusing on different characteristics, such as the depth of the reasoning tree, the need for external information, etc. We focus on the main partition and its subpartitions (i.e., BGQA-1, BGQA-2, BGQA-3), where the number refers to the number of reasoning hops required to answer the question. This dataset also includes annotated chain-of-thoughts (CoT); therefore, we use their annotated input (“example”) as the input prompt and their annotated CoT (“proof”) as the expected output.
We include more details about the datasets in Appendix A, a formal definition of the prompts in Appendix B, and examples in Appendix G.
4.2 LLM Setup.
We perform our study using gpt-35-turbo-0613-16k through the Azure OpenAI service. All of our prompting methods are implemented using the Langchain library. [^2] https://github.com/langchain-ai/langchain We set the decoding temperature to zero and use greedy sampling to make the outputs deterministic. We execute our prompts with in-context learning. However, the amount of demonstrations is constrained by their long length and the context window of the model, especially in code prompts, which are longer than text prompts because they include the original natural language text and the corresponding code. Specifically, we used between three and nine demonstrations except for code prompts in BGQA-2 and BGQA-3 where the maximum possible is 6. In other words, we provided from one to three demonstrations per class. For each experiment, we use a random sample from the training set as demonstrations. A detailed analysis of the impact of the number of demonstrations in the performance of the prompts is provided in Section 5.4. We also report the costs of our prompts in Appendix C.
4.3 Evaluation.
We follow the evaluation metrics used in the original datasets. For CondQA, we report the F1 token overlap between the predicted answer and the label, while for BGQA, we report accuracy. We run the main experiments multiple times with different random seeds: two times on BGQA and three times on CondQA. We report the average and standard deviation performance across these runs. For the subsequent analyses of code prompts, we run each experiment once due to the inference costs.
Prompt Method CondQA BGQA-1 BGQA-2 BGQA-3
Text Prompt $56.32\pm 1.06$ $62.80\pm 0.28$ $55.00\pm 0.56$ $47.59\pm 0.00$
Code Prompt 59.17 $\pm$ 0.74 69.40 $\pm$ 1.13 62.00 $\pm$ 1.70 55.30 $\pm$ 0.42
Atomic Statements 55.94 57.80 47.80 35.60
Code $\rightarrow$ NL 55.42 63.00 55.60 48.00
Table 1: Comparison of text prompt and code prompts on the validation set of each dataset. Last two rows analyse the impact of the input format on text prompts.
5 Experiments
We first compare the performance of the two prompting methods — text prompts and code prompts. We then conduct extensive ablation experiments to determine whether it is the code-translated instances what improve performance or if the improvement is caused by implicit text simplification (Section 5.2) or by models merely being exposed to code within prompts and not necessarily the instances translated to code (Section 5.3). Finally, we show that code prompting is more sample efficient (Section 5.4) when compared to text prompting and that models prompted with code exhibit superior state tracking capabilities (Section 5.5).
5.1 Code Prompting Improves over Text Prompting
Overall, as shown in Table 1, code prompts consistently outperform text prompts. In particular, we can observe a significant performance boost of $7.71$ points on BGQA-3, $7.4$ on BGQA-2, $6.6$ on BGQA-1, and $2.6$ points on CondQA. Furthermore, we can also observe how the gap between both prompting methods increases in BGQA as the difficulty of the task increases. This clearly indicates the potential benefits of using code to elicit reasoning abilities, especially on the multi-hop subsets.
We note that these consistent and substantial gains using code prompts are obtained despite the straightforward transformation of the text prompt, which does not incorporate new information as shown in Figure 2. This suggests that the code format exhibits characteristics that trigger the conditional reasoning abilities of text+code LLMs.
In the following experiments, we perform detailed ablation studies confirming that the performance improvements indeed originate from the code format and render additional insights into the potential reasons behind this phenomenon.
Prompt CondQA CondQA-YN BGQA-1 BGQA-2 BGQA-3
Code Prompt $\mathbf{60.04}$ $\mathbf{78.94}$ $\mathbf{70.2}$ $\mathbf{63.2}$ $\mathbf{55.0}$
Anonymous code $58.42$ $76.02$ $63.6$ $58.4$ $51.0$
Random code $56.64$ $76.27$ $62.8$ $54.0$ $45.2$
- Comments N.A. $64.92$ $53.5$ $47.0$ $49.8$
Table 2: Ablation study on the relevance of code semantics within Code Prompts. CondQA-YN represents the partition of ConditionalQA with yes-no answers.
5.2 Code Format Elicits Reasoning Abilities
We are now interested in whether the observed improvements in code prompting originate from the input instances being formatted as code or if they are caused by the implicit simplification of text that occurs when the instances are translated into code. To investigate this, we devise experiments with prompts that represent the intermediate states between natural language and code.
Inspired by Min et al. (2023), in our first study, we transform each NL sentence [^3] We only transform the facts in BGQA since transforming the rules into atomic statements as well yields worse results. into a sequence of atomic statements, which we then append to the original sentence. In this way, the atomic statements can be seen as defining variables for each key entity in the text. Therefore, this new type of prompt would resemble code but without control flow and in natural language form. For instance, for the sentence Applying for the legal right to deal with someone’s property, money and possessions (their estate) when they die is called applying for probate, some atomic statements would be: i) Applying for the legal right is a process and ii) The process is called ‘applying for probate’. A complete example is provided in Appendix D.
The prompt retains access to the original instance text (i.e., no loss of information) but is also augmented by simplified sentences in the form of atomic statements. This setup allows us to investigate whether the simplicity of the input prompt is the cause of improved reasoning abilities, regardless of the text and code formats.
In our second experiment, we investigate whether the semantics of the code statements and not the form of the code are the reason behind the performance boost. For this purpose, we back-transform the code prompts into NL such that the reasoning statements (i.e., the if conditions) are clearly and concisely stated in natural language. Specifically, we map every variable into the format Key entity: variable without snake case. For instance, the variable husband_pass_away from Figure 2 would be back-transformed as Key entity: husband pass away. To transform the if statements, we create a translation prompt by providing four demonstrations. These demonstrations simply translate the conditional statements within the  code-formatted instance back into natural language in the same manner as the variables. This makes the back-translated text as close as possible to the code text. We provide examples of this in Table 6 from Appendix E.
From the results in Table 1, we see that (1) prompting with atomic statements does not yield performance improvements w.r.t. text prompts, and (2) mapping back from code to NL provokes performance drop w.r.t. code prompts, yielding similar results to the original text prompt. In the first case, we observe that the performance drop w.r.t. the original text prompt can be due to the sampling variance since the results are within the standard deviation of text prompts. However, we obtain a more acute drop for BGQA. This can be attributed to the format of this dataset. The atomic statements method we use (Min et al., 2023) was devised for general texts such as Wikipedia pages. However, BGQA is a logic-based dataset, where the input facts are statements that are already minimal units of information and do not follow the style of general documents. Thus, creating atomic statements from these sentences can break the sentence structure needed to track the attributes of the subjects and objects of the sentences.
These results indicate that the enhanced complex reasoning abilities do not originate from input text simplification nor the conditional statements being extracted from the instance sentences, but rather the code format of the input instances.
5.3 Code Semantic is Important
Previously, we have shown that the code format is necessary to elicit the reasoning abilities of text+code LLMs. Now, we aim to investigate which aspects of code are needed for this. In particular, we evaluate the impact of retaining the natural language text of the original instance within the code comments and the importance of the semantics of the code. To analyze the former, we have removed the code comments that include the original natural language text from the input and evaluated the performance of the new prompts. To analyze the latter, we perturbed the code to anonymize the variables and functions, as well as added random code whose semantics are completely irrelevant to the original natural language text. In the latter two cases, the code comments remain unmodified (examples illustrating them are provided in Table 7 from Appendix E). Since CondQA includes span answers and removing the NL text would make it impossible for the model to generate the span, we only report performance on the yes-no answers partition (i.e., CondQA-YN).
Table 2 shows that removing the NL text in the code comments yields a performance drop of $14.02$ points on CondQA and a performance drop between $16.7$ and $5.2$ on BGQA. This significant and consistent decrease in all datasets confirms that retaining NL text in comments is vital for the LLM to understand the input problem.
Code perturbations (anonymous code and random code) also confirm the importance of code semantics to elicit reasoning abilities. When we use anonymized code, we observe a performance reduction of almost $2$ points on CondQA and a decrease between $6.6$ and $4$ in BGQA. The decrease is even larger when the code is randomized, with drops of more than $3$ points on CondQA and between $7.4$ and $9.8$ on BGQA. This more pronounced drop is expected since the semantics and logic of the code mismatch the NL text, whereas anonymous code maintains the same logic on both NL and code.
Furthermore, we also observe that the performance of random code prompts is similar to that of text prompts (Table 1) in all datasets. This can be interpreted as the model being able to identify the irrelevance of the code to the text. Hence, the model disregards the code to solely focus on the code comments (i.e., the natural language text). This could be possible thanks to the provided demonstrations, which show answers that only refer to the natural language text.
These results confirm that code alone does not trigger reasoning abilities, and instead, the combination of code that represents the original natural language instance and the NL text exploits the potential of LLMs.
5.4 Code Prompts are More Sample-Efficient at Eliciting Reasoning Abilities
Given our observations that code prompts trigger conditional reasoning abilities better than text prompts, it is natural to ask the follow-up question: are code prompts also more sample-efficient at triggering conditional reasoning abilities when compared to text prompts? To answer this, we evaluate how the overall performance of the model changes with respect to the number of demonstrations.
In the following experiments, we compare the performance gap of the two prompting methods under different numbers of demonstrations. Figure 3 shows that when we only provide one demonstration per class (i.e., answer type in our datasets), the performance gap is the largest across all datasets. As expected, this gap decreases when we provide more demonstrations. These results indicate that code prompts trigger conditional reasoning more efficiently than text prompts, and this is one of the reasons for its superior performance. This phenomenon has further applicability in long-document and very low resource settings, where only one demonstration is available or fits within the input window, and therefore, it would be advisable to convert that demonstration into code.
Figure 3: Performance comparison between text prompts (blue) and code prompts (green) using 1, 2, and 3 demonstrations per class.
5.5 Code Prompts Improve Variable Tracking in LLMs
Correct Instances Incorrect Instances
Dataset Text Prompts Code Prompts Text Prompts Code Prompts
CondQA 71.08 4.39 60.79 11.39
BGQA-1 39.33 8.84 51.65 22.12
BGQA-2 44.79 15.04 52.54 24.75
BGQA-3 54.01 14.21 52.13 16.98
Table 3: Comparison of the percentage of memory errors made by GPT 3.5. For each dataset, we separately compute memory errors for the instances where the model gives the correct and incorrect answers. Lower is better.
We hypothesize that one of the reasons for the superior performance of code prompting is an improved ability to track the states of variables or key concepts. This hypothesis is based on the intuition that, for natural language in general, local context is the most important part of generating the next token Khandelwal et al. (2018); Sun et al. (2021). However, generating code is often more challenging because code frequently refers to previously defined functions, which can be dozens or even hundreds of lines apart. This resembles multi-hop reasoning, where the model may need to reference a key entity dozens of lines before. Therefore, an improved ability to look for distant co-references can be beneficial to multi-hop reasoning, which is also needed to solve our datasets.
[…]
We acknowledge that the output text may not be completely faithful to the internal beliefs of the model (Lyu et al., 2023). Therefore, we first test whether this experiment can be a proxy metric of the internal belief of the model. To do this, we compare the memory error percentage of the prompting methods in instances where the model solves (i.e., correct instances) and does not solve (i.e., incorrect instances) the question. If incorrect instances yield a higher memory error, this would indicate that the model struggles more to remember the state of those variables, and this would make it fail to conduct the reasoning process. Therefore, our probes would be a proxy metric of the internal belief of the model.
Table 3 shows the results of this comparison. We observe that all prompting methods in all datasets consistently make more memory mistakes on incorrect instances than on correct instances, with the exception of text prompts on CondQA. However, the memory error in this case is significantly high, which may suggest that the model is not able to track entities correctly in any case. Therefore, we can use this experiment as a proxy measure of the memory of the model.
From the same table, we can observe that Text Prompts make significantly more memory errors than code prompts on all datasets. Specifically, the gap is consistently more than 30% with peaks on CondQA (66.69%) and BGQA-3 (39.8%). Therefore, this experiment empirically confirms our hypothesis that code prompts can track the state of the variables/key entities better than text prompts.
5.6 Qualitative Analysis
All our previous experiments have focused on the accuracy of the answers. Obtaining a correct answer is correlated to a correct reasoning process. However, it is possible to obtain a correct answer despite an incorrect reasoning chain. Due to the difficulty of automatically evaluating the chain of thoughts, we perform a limited human evaluation. We sample ten instances from BGQA-3 (i.e., the dataset with the most complex reasoning chains) where both text and code prompts return the correct answer and manually inspect the quality of the reasoning chains. Based on this limited sample, we observe that text prompts conduct a 100% correct reasoning chain in only two cases, while code prompts do so in three cases. We identify two main types of errors: (1) wrong conditional reasoning and (2) commonsense errors. We provide examples of those in Table 4. This observation raises the question of whether the generated chain of thought is not faithful to the internal reasoning of the model, as suggested by (Lyu et al., 2023) or whether the model generated the right answer from a greedy attempt to reach the closest plausible conclusion (code prompts shows a reasoning error in the last step of the CoT in three out of seven cases). We leave the analysis of faithfulness of chains of thoughts as future work.
Error Type Example Explanation
Commonsense
We know the catfish has a harmonica, and according to Rule3 "if the catfish has something to sit on, …
The model believes a harmonica is something to sit on.
Conditional Reasoning
# We know the lion does not remove from the board one of the pieces of the dog, and according to Rule5 "if something becomes an enemy of the squid but does not remove from the board one of the pieces of the dog, then it steals five points from the mosquito."
become_enemy(lion, squid)==True
not remove_piece(lion, dog)==True
steal_points(lion,mosquito,5)=rule5(lion)
steal_points(lion, mosquito, 5)==True
We do not know  become_enemy(lion, squid) == True, but if we assume this, we reach the question variable, so we get an answer to the question.
Table 4: Examples of reasoning errors on correct instances by text and code prompts. Underline text is the cause of the error.
6 Conclusions and Future Work
This work demonstrates that conditional reasoning abilities can be triggered by using code prompts in large language models (LLMs) of text and code. These code prompts contain the original natural language (NL) formulation as a code comment and code that formulates the logic of the text (e.g., documents and questions). To create these code prompts, we use in-context learning to teach an LLM how to automatically conduct such a transformation. Through multiple experiments, we show that code prompts trigger conditional reasoning abilities. They consistently and substantially outperform text prompts across two conditional reasoning datasets by between $2.6$ and $7.7$ points. Our experiments show that even simple code can be beneficial as long as it closely follows the semantics of the NL text and is accompanied by the original NL text. In addition, code prompts require fewer demonstrations, making them more efficient than vanilla text prompts. Our experiments also suggest that code prompts achieve the performance boost due to their superior ability to track the state of variables or key entities.
In our future work, we plan to extend our experiments to other text+code LLMs. We also plan to conduct experiments investigating how pretraining on text, code, and text+code affects the triggering of reasoning abilities. Lastly, we would like to extend our experiments to a wider range of reasoning abilities to verify the generalization of our findings.
Limitations
Transforming a natural language problem into code requires an intermediate step that raises the cost of the whole pipeline. However, this mapping is not a complicated task. Therefore, we believe it would be possible to train a small generative model to do it instead of using a large language model. In this way, we could minimize the cost of using code prompts without affecting its performance.
Due to the costs of very large language models and the large size of BoardgameQA, we only ran the experiments two times with different random seeds. However, the variance is small enough to ensure the representativeness of the results.
Acknowledgements
This work has been funded by the German Research Foundation (DFG) as part of the UKP-SQuARE project (grant GU 798/29-1), by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE, and by the European Union (ERC, InterText, 101054961). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research).
Lastly, we thank Max Glockner, Jonathan Tonglet, and Sheng Lu for their insightful comments and suggestions on a draft of this paper.
Appendix A Datasets
In ConditionalQA, we use an oracle retriever to retrieve the relevant passages to the question. This retriever is based on the sentences annotated as evidence for the answer (i.e., rationales). We concatenate all sections that include one rationale and use the resulting passage as input document.
The sizes of all the datasets used in this work are provided in Table 5.
Dataset Training Size Dev Size
CondQA 2338 285
BGQA-1 1000 500
BGQA-2 1000 500
BGQA-3 1000 500
Table 5: Sizes of the datasets.
Appendix B Prompt Formulation
CONDQA.
Firstly, we define the different components of a data point: scenario ($S$), question ($Q$), document ($D$), rationales ($R$), and answer ($A$). Then, the text prompt $tp$ is defined as follows:
$\begin{split}tp=\text{"Question:"}+S+Q+\text{"Document:"}+D\\ +\text{"Let's think step by step"}\end{split}$ (1)
where $+$ represents the string concatenation operator. Then, the output format, $to$ is:
$to=R+\text{"Answer:"}+A$ (2)
For code prompts, we first define a function ${C:\mathbb{NL}\rightarrow\mathbb{C}}$ that maps a natural language text into code as shown in Figure 2. Then, we define code prompt $cp$ as follows:
$\begin{split}cp=\text{"\#Question:"}+C(S)+C(Q)+\\ \text{"\#Document:"}+C(D)\\ +\text{"\#Let's think step by step"}\end{split}$ (3)
Similarly, we define the output format, $co$, as:
$co=R+\text{"\#Answer:"}+A$ (4)
BGQA.
Firstly, we define the components of a data point in this dataset: facts ($F$), rules ($R$), and questions ($Q$). Therefore, our text prompt is defined as follows [^4] BGQA provides a field example with all the variables of the dataset concatenated with descriptions. We use this field as text prompt.:
$\begin{split}tp=F+R+Q\end{split}$ (5)
This dataset also provides the CoT that leads to the answer. Therefore, we use that CoT as the expected output.
For code prompts, we follow the same approach as with the previous dataset. We define code prompts, $cp$, as follows:
$tp=C(F)+C(R)+C(Q)$ (6)
with the output format ($co$) being:
$co=C(cot)$ (7)
Appendix C Costs
Running a data instance from ConditionalQA with code prompts costs \$0.04 while with text prompts \$0.01. On BoardgameQA-depth 3 (i.e., the partition with the most expensive prompts), the costs per question are \$0.02 and \$0.03 for text and code prompts, respectively.
Appendix D Atomic Statements
Original sentence: <p>Applying for the legal right to deal with someone’s property, money and possessions (their estate) when they die is called applying for probate.</p> Atomic statements: Applying for the legal right is a process. The process is called ’applying for probate’. The legal right is to deal with someone’s property, money, and possessions. The someone is a person who has died. The property, money, and possessions are collectively called the ’estate’.
Appendix E Examples of Code Ablations
An example of a back-translated code into natural language is provided in Table 6. We can observe in both examples that the resulting natural language (NL) text is extremely similar to the original code. In addition, in the second example (BGQA), Rule2 is much simpler after the back-translation than its original description in NL.
Table 7 shows examples of the multiple code ablations we conducted in Section 5.3. Random code replaces the code with a piece of code from another data point. In this way, the semantics of the text and code mismatch while we keep the code syntactically correct.
Appendix F Variable Tracking Setup
Extracting key entities in BoardgameQA.
This dataset provides a list of “facts,” which are short and concise sentences describing the state of a key entity. Therefore, we use them without alterations as the key entities to ask for.
Extracting key entities in ConditionalQA.
This dataset provides a scenario describing the background information of the person posing the answer. Since this scenario is a free-form text, we follow (Min et al., 2023) to extract atomic statements and use them as the key entities to ask for.
Code Prompting variables
. To probe the variable tracking abilities of code prompts, we use the variables defined in the “facts” and “scenario” of BoardgameQA and ConditionalQA, respectively.
Probing memory at different steps in the Chain-of-Thought.
[…]
Number of probes.
For each dataset instance, we run $num\_facts\times num_{s}teps_{c}ot$ probes, which makes this experiment very costly. Thus, we aim to maximize the number of instances probed while keeping the costs down. To do so, we use a sample of 50 instances for each dataset partition of BoardgameQA, except for Board3, where we used 20 instances ($\approx 700$ probes) because of the cost of the experiment. Due to the length of the demonstrations of ConditionalQA and its impact on the costs, we sample five facts and three partial CoTs for each instance, yielding an upper-bound of 15 probes per instance, and run the probes for 30 instances for each dataset partition (i.e., correct and incorrect instances).
Prompt Probes.
In all cases, we follow the following format: Sys. Prompt; ICL Demonstrations; Input Instance; Partial CoT; Probe.
The probe for text and code prompts in BoardgameQA is: “Now, I want to ask you about the value of some key entities you used. Your answers must be ‘yes‘, ‘no‘, or ‘unknown‘. It is very important that you only write one word. Is it true that {fact}?”
The probe for text prompts in ConditionalQA is: “Now, I want to ask you about the value of some key entities you used. Your answers must be “True”, “False”, “unknown”, or a string. It is very important that you only write the exact value. From the speaker perspective, is it true that {fact}?”
The probe for code prompts in ConditionalQA is: “Now, I want to ask you about the value of some key entities you used. Your answers must be “True”, “False”, “unknown”, or a string. It is very important that you only write the exact value. What is the value of the variable {var}?” A real example is provided in Table 8.
Type Text
Code
# <p>You can apply to become the estate’s administrator if you are 18 or over and you are the most ‘entitled’ inheritor of the deceased’s estate. This is usually the deceased’s closest living relative.</p>
if applicant_age >= 18 and entitled_inheritor and closest_relative:can_apply_estate_administrator = True
Code $\rightarrow$ NL
<p>You can apply to become the estate’s administrator if you are 18 or over and you are the most ‘entitled’ inheritor of the deceased’s estate. This is usually the deceased’s closest living relative.</p>
if you are 18 or over and you are the most entitled inheritor of the deceased’s estate and you are the closest living relative, you can apply to become the estate’s administrator
Code
# Rule2: Be careful when something removes from the board one of the pieces of the dog and also becomes an enemy of the catfish because in this case it will surely not burn the warehouse of the mosquito (this may or may not be problematic)
rule2(something) = remove(something, piece_of(dog)) & enemy(something, catfish) => not burn(something, warehouse_of(mosquito))
Code $\rightarrow$ NL
Rule2: If something removes from the board one of the pieces of the dog and also becomes an enemy of the catfish, then it does not burn the warehouse of the mosquito
Table 6: Example of a back-translation ${\mathbb{NL}\rightarrow\mathbb{C}}$ in ConditionalQA and BGQA-3. Text in bold represents the main modification.
Type Text
Original Code
# <p>To be eligible you must have left your country and be unable to go back because you fear persecution.</p>
if left_country_and_fear_persecution:
eligible_for_asylum = True
Anonymous Code
# <p>To be eligible you must have left your country and be unable to go back because you fear persecution.</p>
if var_1
var_2 = True
Random Code
# <p>To be eligible you must have left your country and be unable to go back because you fear persecution.</p>
if value_of_property_gone_down_by_more_than_50:
eligible_to_claim = True
getting_housing_benefit = True
Table 7: Examples code ablations.
Section Role Message
Problem instance
Human
Question: My brother and his wife are in prison for carrying out a large fraud scheme. Their 7 and 8 year old children have been living with me for the last 4 years. I want to become their Special Guardian to look after them permanently. How long will it be before I hear back from the court?
Document: <h1>What is a special guardian</h1> <p>You can apply to be a child’s special guardian when they cannot live with their birth parents and adoption is not right for them.</p> …
Answers can be "yes" or "no". Let’s think step by step:
Partial CoT
AI
[…]
Probe
Human
Now, I want to ask you about the value of some key entities you used. Your answers must be ‘True‘, ‘False‘, ‘unknown‘, or a string. It is very important that you only write the exact value. From the speaker perspective, is it true that the children have been living with me for the last 4 years?
Probe
AI
True
Table 8: Variable Tracking Example. Underlined text represents the variable to probe. Partial CoT is not the complete answer. The generation was stopped, and only the first step was used in this probe.
Appendix G Prompt Examples
System: You are a helpful assistant that answers questions given a document. Answers must be a short span of the document. You have to extract the span from the document. Do not write anything else. I will give you some examples first.
ICL Demonstrations…
Human: Question: My brother and his wife are in prison for carrying out a large fraud scheme. Their 7 and 8 year old children have been living with me for the last 4 years. I want to become their Special Guardian to look after them permanently. How long will it be before I hear back from the court?
Document: <h1>What is a special guardian</h1>
<p>You can apply to be a child’s special guardian when they cannot live with their birth parents and adoption is not right for them.</p>
<p>You’ll be responsible for looking after the child until they’re 18 (unless the court takes your responsibility away earlier).</p>
<p>You’ll make all day to day decisions about the child, for example schooling and medical treatment. You do not have to discuss these decisions with the birth parents.</p>
<p>You’ll need to get the consent of everyone who has parental responsibility for the child before you make some important decisions, for example:</p>
<li>changing the child’s surname</li>
<li>putting the child up for adoption</li>
<li>taking the child abroad for more than 3 months</li>
<li>the child having surgery for reasons other than improving health, such as circumcision, sterilisation or cosmetic surgery</li>
<p>If you cannot get consent, you can ask the court to decide. Use the form ‘Make an application in existing court proceedings related to children’ (form C2).</p>
<h1>After you apply</h1>
<p>Within 10 days of receiving your application the court will send you a case number and a date for a meeting to set out:</p>
<li>a timetable for your case</li>
<li>how it will be dealt with</li>
<p>This meeting is called a ‘first directions hearing’.</p>
<p>You must go to all hearings you’re told to unless the court excuses you. If you’re not able to go, contact the court office.</p> Answers must be a short span of the document. You have to extract the span from the document. Do not write anything else. Let’s think step by step:
Table 9: Text prompt Example for ConditionalQA
System: You are a helpful assistant. Your task is to process a pseudo-code that describes a question and a document. You need to reason using that document and the comments to return the answers. Answers must be a short span of the document. You have to extract the span from the code comments. Do not write anything else. I will give you some examples first.
ICL Demonstrations…
Human: # Question: My brother and his wife are in prison for carrying out a large fraud scheme. Their 7 and 8 year old children have been living with me for the last 4 years. I want to become their Special Guardian to look after them permanently. How long will it be before I hear back from the court?
maximum_redundancy_pay = 16320
housing_standards_and_procedures_in_Northern_Ireland = True
ensure_vehicle_taxed_in_UK = True immigration_advisers_can_help_with_representation_at_tribunal = True
supply_protective_clothing_and_equipment = True
CBT_required_for_moped_and_motorcycle = True
court_response_time = None # This is the variable that answers the question
# <h1>What is a special guardian</h1>
# <p>You can apply to be a child’s special guardian when they cannot live with their birth parents and adoption is not right for them.</p>
if attorneys_appointed_jointly:
all_attorneys_must_agree_to_make_decision = True
disability_or_severe_disability_element_of_working_tax_credit = True
mugging_without_physical_harm_emergency = True
# <p>You’ll be responsible for looking after the child until they’re 18 (unless the court takes your responsibility away earlier).</p>
work_temporarily_for_hirer = True
# <p>You’ll make all day to day decisions about the child, for example schooling and medical treatment. You do not have to discuss these decisions with the birth parents.</p>
accounts_and_tax_returns_cover_financial_year = "1 June to 31 May"
employer_operating_PAYE = True
# <p>You’ll need to get the consent of everyone who has parental responsibility for the child before you make some important decisions, for example:</p>
# <li>changing the child’s surname</li>
# <li>putting the child up for adoption</li>
# <li>taking the child abroad for more than 3 months</li>
# <li>the child having surgery for reasons other than improving health, such as circumcision, sterilisation or cosmetic surgery</li>
managed_by_fit_and_proper_persons = True
check_court_order_for_authorization = True
considering_fostering = True
if not_connected_to_mains_sewer:
septic_tank_used = True
can_claim_tax_relief_if_taxed_twice = True
extra_support_for_disability = True
if operator_of_septic_tank_or_treatment_plant:
follow_general_binding_rules = True
# <p>If you cannot get consent, you can ask the court to decide. Use the form ‘Make an application in existing court proceedings related to children’ (form C2).</p>
appeals_decision_time = "several months"
if worker and informal_resolution_not_satisfactory:
formal_grievance_complaint_possible = True
time_limit_for_backdating_claims_services = 6
# <h1>After you apply</h1>
# <p>Within 10 days of receiving your application the court will send you a case number and a date for a meeting to set out:</p>
# <li>a timetable for your case</li>
# <li>how it will be dealt with</li>
# <p>This meeting is called a ‘first directions hearing’.</p>
committee_recommendations_go_to_Prime_Minister = True
check_adviser_registration = True
meet_manning_levels = True
recognised_as_charity_or_CASC = True
apply_for_visa_for_other_reasons = True
debt_paid_off = True
if special_educational_needs_and_disabilities:
affects_behaviour_or_socialisation = True
# <p>You must go to all hearings you’re told to unless the court excuses you. If you’re not able to go, contact the court office.</p>
payslip_can_include_tax_code = True
VAT_zero_rate = 0
gas_equipment_installed_and_maintained_by_Gas_Safe_registered_engineer = True
# Question: My brother and his wife are in prison for carrying out a large fraud scheme. Their 7 and 8 year old children have been living with me for the last 4 years. I want to become their Special Guardian to look after them permanently. How long will it be before I hear back from the court?
# Answers must be a short span of the document. You have to extract the span from the code comments. Do not write anything else.
# Let’s think step by step:
Table 10: Code Prompt Example for ConditionalQA
System: You are a question-answering system that solves the problem of reasoning with contradictory information guided by preferences over sources of information. You must explain your answers step by step.
ICL Demonstrations …
Human: A few players are playing a boardgame
The current state of the game is as follows
The amberjack struggles to find food
And the rules of the game are as follows
Rule1: If the amberjack has difficulty to find food, then the amberjack removes from the board one of the pieces of the carp
Based on the game state and the rules and preferences, does the amberjack remove from the board one of the pieces of the carp?
AI:
Table 11: Text prompt Example for BGQA-1
System: You are a large language model of code that can interpret code. You are given a pseudo-code that resembles to first-order logic that models some scenario. You will be given a question and you have to answer it step by step. You can use a rule if and only if you know the antecedent of the rule.
ICL Demonstrations
Human: # A few players are playing a boardgame
# The rules of the game are as follows
# Rule1: If the amberjack has difficulty to find food, then the amberjack removes from the board one of the pieces of the carp.
rule1() = difficulty_finding_food(amberjack) => remove_piece(amberjack, carp)
# The current state of the game is as follows
# The amberjack struggles to find food.
difficulty_finding_food(amberjack) = True
# Based on the game state and the rules and preferences, does the amberjack remove from the board one of the pieces of the carp?
question = remove_piece(amberjack, carp)
AI:
Table 12: Code prompt Example for BGQA-1